NAOJ GW Elog Logbook 3.2

I placed a lamp at a distance D from the convex surface and I looked
for the image reflected by the second face of the mirror (which is flat).
The image is real and placed at a distance D' from the convex surface.
I changed D until D'=D. In this situation the focal length f is
f = D = D' = 1.305 +/- 0.01 m
I can deduce the ROC from the expression ROC = f * (n-1) where n is the
fused silica index of refraction in the visible (green) i.e n=1.4607 at
lambda = 532 nm. I found ROC=0.601 +/- 0.005 m.
In conclusion, within the error of the measurement the mirror ROC
is equal to the required value (which is 0.6 m).

This morning I have removed the supports used to glue the standoffs on the BS telescope mirror that I glued last wednesday evening. For the moment, they look well attached to the mirror.
The magnets were glued on tuesday and their supports were removed on wednesday in order to rotate the mirror and allow for the glueing on the standoffs.
Since BS mirror has a different disposition of magnets with respect to test masses, the glueing has to be done in 2 times.
http://www2.nao.ac.jp/~gw-elog/osl/uploads/260_20160627054235_mirrormagnet.jpg
Both magnets and standoff have been glued using a new set of MasterBond EP30-2, stored in the fridge of TAMA.

In the optical lever scheme, a beam is sent to the center of the the mirror with an angle α.
A rotation of the mirror around its vertical axis (a yaw) of an angle β, corresponds to an angular displacement of the reflected beam of 2β.
This produces a shift of the beam's position on a PSD placed at a distance L from the mirror of
X = L tan (2β) ≈ 2Lβ
Now, if the beam impinges (with an angle α) at a distance d from the center, the shift on the PSD becomes
Xd = L tan (2β) + d (sinβ/cos(α+β)) ≈ 2Lβ + d β (1/(cosα - βsinα))
In the case of the beamsplitter, the optical lever beam is currently sent with a very small incidence angle (α ≈ 0) and impinges at a distance d (in the orizontal direction) of about 3.5 cm from the mirror center. The distance between the mirror and the PSD (arm of the optical lever) is about 60 cm.
The relative errore due to the mis-centering is (Xd-X) / X = d/(2L) = 3%
which is small. So it seems not so crucial to hit the center of the mirror.
For what concers the pitch, I would say that as long as the beam is on the orizontal axis (as in our case), its mesurement is not affected by the mis-centering.

I disassembled one of the two HEPA filter of the cleanbooth for the absorption measurement system ( Model MAC-103 ) in order to check the conditions and to figure out whether and how to change it. The circular prefilter is very easy to replace and very dirty, I think we should change it. The real HEPA filter inside is quite dirty but is not easy to replace because the filter is kind of glued at the frame, so maybe we can contact the company. I hope there is the possibility to change only the filter, and not the whole fan.
The filter's fan part was fine, so it was enough to buy only the filter part. New HEPA filters were delivered to Tama last friday. I washed the prefilters with water, I cleaned the fans, replaced the filters and placed them back to the top of the absorption bench clean booth.

And we could not kill one higher order mode to get better mode matching, and also there is one mode jump out from time to time when we try to align the cavity. The final efficiency we got is about 20%(260mW goes into the cavity and 55mW green received at the transmission of the dichoric, 12mW received by the reflection of the dichoric)

Then we increased the power a little bit, do the alignment again, then the error signal and the infrared signal shows in the third picture. There is a flashing spot in the infrared signal.
But the changing part is when we lock the cavity, the 500mW situation has the oscillation also(picture 4), but the second seams not.(pic 5)
According to this circumstance, we at least got one conclusion is that the high frequency oscillation has nothing to do with the green.
Then we rethought the polarization again, found out no matter s or p polarization reflect back, it is should stopped by the Faraday Isolator, but now it seems not.
We decided to check the polarization again from the beginning. Firstly one thing for sure is that, the input infrared beam to the cavity is in p polarization, and the reflect infrared beam is also p polarization.
Then we try to put another FI and half wave plate after the BS mirror and before the telescope to cancel the reflect infrared beam from the cavity. We use the same FI as we used before, but found out the two polarization cube point to different direction of the original one. Then the new put FI did reflect out the reflection beam, and get rid of the oscillation.
Then we guess that the original FI did not do its job maybe because the two wave plate before it did not clean the polarization as they should do.So we put a PBS after the first quarter wave plate and second wave plate, found out exactly what we suspected, so we adjusted them firstly then the Faraday and the other two half wave plate, got pure p polarization inject into the cavity.
Now the oscillation seems not there but the efficiency of green beam is only about 10%, so we tried to increased the alignment, but it seems we should do it from the start. After we make a progress in the alignment, we can check again if the oscillation really disappear. Then we can be sure that it is caused by the polarization problem.

After doing this, we remove all the tall mounts for infrared and do the alignment again both in the infrared and green path. During the alignment, we found out maybe the cavity itself is a little bit tilt up, with one mirror we recover the beam to the right height.
Then we inject both the green and infrared beam into the chamber.
Firstly, adjust the position of G1 mirror in the picture try to get the green beam in the middle of it in horizontal direction. Then use adjust it slightly to get the reflection beam goes to the GR mirror.
Then we adjust the infrared beam, use the last two mirror on the bench in infrared path to adjust the infrared go through the Faraday Isolator, reflect by the R2 mirror and also goes to the GR mirror.
Then after the GR mirror we try to do our best to overlap the infrared and green beam and get a good result, next step is let these two beams goes into the BS chamber and try to overlap them there.
Then we installed the RP1 mirror, this is a mirror with high reflection with 0 zero degree. Then also install the RP2 mirror, this is the mirror has high reflection in P polarization. With this two mirror we got the p polarization reflection of the infrared beam and it arrives about the right position in the bench.

This morning I have removed the support of the standoff glued on saturday. This time it has remained attached to the mirror and it looks ok.

This morning I went to check the status of the magnets and stand-off we glued on thursday on the telescope mirror.
While removing the jigs, a stand-off came off. We probably put to much glue on it and it overflowed a bit on the support causing the stand-off to remain attached to the support.
After removing it, I have cleaned the mirror and the support and glued it again paying more attention to avoid overflow.
Matteo gently touched the other 3 stand-offs and they seem to be well glued. We did not touch the magnets but they also look good.
Some pictures taken by Matteo are attached.
A special thank to Yuefan who woke up at 10 a.m. to prepare the glue!

Today we checked the picomotors in the BS chamber.
BS suspension picomotors
The out-of-vacuum connector is connected to the flange named PO 4 (see first attached picture)
There are two cables (with 6 pins) coming out from the connector, that I labelled BS ALIGNMENT and BS LENGTH.
Here what we found:
BS ALIGNEMENT
- A: PITCH
- B: YAW
- C:
BS LENGTH
- A: Z
- B: X
- C: should be Y but doesn't work!
We checked the connections and they seem to work, so it is probably the picomotor itself which doesn't work although the screw can be rotate by hand.
2'' telescope mirror picomotors
The not-suspended 2" telescope mirror will be equipped with 3 picomotors (pitch, yaw and a translation). See picture 2.
I have checked that the 3 picomotors can work and I prepared a cable for the out-of-vacuum connection.
I re-used a former cable used for PO1 in the PR chamber adding 2 pins for controlling the translation picomotor. The pinout of the modified, cable labelled TELESCOPE 2", is shown in figure 3.

We checkd the ROC of the 4" telescope mirror. To do that we used the mirror to focus the
image of a lamp placed at distance D1 and measured the distance D2 of the image from the mirror.
When the two distances are made equal D1 = D2 = 6.05 m = ROC of mirror.
The error of the measurement is about +/- 2 cm.
The specification was 6 m +/- 0.3 m. Looks good.

Wednesday, March 15th, 2017
We opened the top of the BS vacuum chamber.
Apart from the BS suspension and the corresponding cables for the coils and the
picomotors, there is no additional cable in the vacuum chamber.
There are two masses that were used as counter masses to balance the PO suspensions.

Tuesday, March 14th, 2017
Morning
We close the MC vacuum chamber.
We move the clean-booth around the PRM vacuum chamber.
We opened the PRM vacuum chamber. Then we installed the clean-booth ceiling and run the clean air flux.
We clean the vacuum chamber as well as the floor around the PRM vacuum chamber.
Afternoon
We installed the Faraday-Isolator inside the PRM vacuum chamber.
While doing that, we moved the two steering mirrors equipped with picomotors from
the Faraday Isolator platform on two independent posts so to place them at the right positions
(drawing and pictures will be added later).
The mirrors are mounted on 7" tall posts so to have the mirrors center at the
correct height i.e. 120 cm from the ground and 21.3 mm from the optical bench which
is the same height as the suspended mirror.

Monday, March 13th, 2017
Morning
We moved the PRM clean-booth around the MC chamber.
We open the top of the MC vacuum chamber and protected the inside of the
vacuum chamber with a plastic sheet.
We then install the clean-booth ceiling and operate the air flux.
Afternoon
We remove the Faraday-Isolator as well as the two steering mirrors from the MC chamber.
The two steering mirrors are equipped with picomotors.
Together with the Faraday-Isolator block we removed the corresponding four
beam dampers (pictures will be added later).
We left the cable used to drive the picomotors inside the MC vacuum chamber.

The second one is the green path on the bench now.
The third one is the infrared path, we increase the beam height.

1.There is some high frequency oscillation when the alignment is better, efficiency reach around 40%(Output power of the laser is 500mW). Even if we reduce the power to only 50mW output from the laser, the oscillation still there. We try to change the temperature and increase the power, find out the oscillation disappear, but under this situation, the alignment is not very good, so we need to find out which is the reason that make the oscillation disappear.
2. When we change the PZT output with hand to get near the resonance from one side, the high frequency oscillation will increase first, but then it starts to reduce, when we reach the resonance, we are not at the maximum of the oscillation.
2.We find out there are two TEM00 mode, one is more powerful then the other.
Now we can lock the cavity with 500 gain, and it is stable.
We installed the simple version of the green path, without the EOM, AOM and MZ, just to send the green beam to the chamber to check the position, and for the infrared beam which should go together into the chamber with the green, we increase it to the right height( the height we will reach after we increase the table leg) and align it to let it go through the Faraday Isolator and reflected by a mirror, goes to the direction to the BS chamber.

Today we checked the picotmotors and the cabling of the 2 steering mirrors in the PR chamber. They both work fine.
The connections has been done in order to have two cables exiting from the flange, labelled SM PITCHS and SM YAWS. (See first attached picutre)
According to the hand-pad convention (A, B, C), the connections are the following:
SM PITCHES:
- A: Pitch SM1
- B: Pitch SM2
- C:
SM YAWS
- A: Yaw SM1
- B: Yaw SM2
- C:
Where SM1 is the steering mirror closer to the suspension (on the path of the green beam) and SM2 is the mirrror further from the suspension (on the path of the IR beam). See second attached picture.
NB: cables for the steering mirror picomotors are connected to the flange closer to the optical table( which has also the coils cables), those for the control of the suspension are connected to the opposite one.

I used the SolidSpec3700 spectrophotometer at ATC to measure the transmission spectrum of the Crystalline coating from CMS.
We have 2 samples, one is AlGaAs transferred on a silica substrate and the other one is transferred on a sapphire substrate.
The settings of the instrument are:
- Slit width = 2nm
- Sampling = 0.5nm
- Range 1400nm - 600nm
- Lamp switch point: 840nm
There are some little differences between the two samples:
- they are shifted by about 3nm
- silica substrate sample has larger side oscillations

Today we checked the picomotors of the suspension in the PR chamber (a part of the filter cavity telescope).
The suspesion is equipped with 5 picomotors and 2 piezo.
Starting from the top (see attached picture):
1° stage:
- YAW (orizontal) + piezo
- PITCH (vertical) + piezo
2° stage:
- Y (defined as the vertical translation of the mirror)
3° stage:
- Z (defined as traslation along the beam axis ( perpendicular to the mirror))
- X (the left one)
In order to control the picomotor we have used as out-of vacuum cable the original cable used for the double pendulum suspension in NM1.
Pinout: http://tamago.mtk.nao.ac.jp/tama/ifo/www_suspension_install_status/burndy.jpg (we are not sure that our definitions of X, Y, Z agree with that reported in this link)
There are two cables (with 6 pins) coming out from the connector, repectively named EF ALIGNMENT and EF LENGTH. We have connected them to the picomotor driver and checked the degree of freedom controlled by each of them.
NB: each cable can move 3 picomotors, which using the hand-pad convention are named A, B, C.
Here what we found:
EF ALIGNEMENT
- A: PITCH
- B: YAW
- C:
EF LENGTH
- A: Z
- B: X
- C: Y
We checked that all the picomotors work in both directions.

- The pressure was 1.2 Torr before vent.
- The GV in the mid point is opened.
- The TMP and the Drypump can work.